# Solving the Inverse Problem of Electrocardiography for Cardiac Digital Twins: A Survey

### v2

@@ Abstract
Cardiac digital twins are personalized virtual representations used to understand complex heart mechanisms. Solving the ECG inverse problem is crucial for accurate virtual heart modelling, enabling the derivation of internal electrical activity information from recorded surface potentials. Despite challenges from cardiac complexity, noisy ECG data, and computational efficiency, recent advancements hold significant promise for enhancing virtual heart modelling, ultimately advancing precision medicine in cardiology. This paper aims to provide a comprehensive review of the methods of solving ECG inverse problem, the validation strategies, the clinical applications, and future perspectives. For the computing methodologies, we broadly classify state-of-the-art approaches into two categories: deterministic and probabilistic methods, including conventional and deep learning-based techniques. Integrating physics laws with deep learning models holds promise, but challenges such as capturing dynamic electrophysiology accurately, accessing accurate domain knowledge, and quantifying prediction uncertainty persist. Integrating models into clinical workflows while ensuring interpretability and usability for healthcare professionals is essential. Overcoming these challenges will drive further research in cardiac digital twins.
@@
Cardiac electrophysiology, inverse problem, computational model, cardiac digital twin, survey.
@@ 1 Introduction
Electrocardiogram (ECG) is a non-invasive technique used to measure the electrical activity of the heart and identify potential abnormalities [1]. It consists of a number of electrodes placed on the skin of the patient, typically on the chest, arms, and legs. These electrodes record the changes in electrical potential on the body surface, which are generated by the underlying electrical activity of the heart. Virtual modelling of cardiac electrophysiology (EP) for digital twins has shown great promise in numerous clinical applications, such as the stratification of arrhythmia risk [2] and the ablation guidance of persistent atrial fibrillation (AF) [3]. However, obtaining numerical solutions to the governing partial differential equations (PDEs) in cardiac EP (ECG forward problem) is quite challenging. Several forward physical modelling approaches have been employed, such as the bidomain model, the monodomain model, and the Eikonal equation. The biodomain model provides information about the ionic currents [4], the monodomain model describes the electrical propagation [5], and the Eikonal equation models wavefront propagation [6]. To obtain a complete understanding of the electrical activity of the heart, we need to estimate the activity inside the heart from the measured ECG signals, namely ECG inverse problem [7]. Fig. 1 presents the inverse inference procedure for reconstructing the cardiac activity from ECG data. By solving the inverse problem, one can obtain valuable information about the function of the heart, which can aid in the diagnosis and monitoring of various cardiac conditions, as well as the development of personalized treatments for cardiac disorders.
However, directly predicting cardiac activity from standard ECG signals (often recorded using a limited number of electrodes, such as the 12-lead ECG with ten electrodes) is a challenging task. This is mainly because the limited number of electrodes in standard ECGs makes it difficult to capture the fine-grained details of cardiac electrodynamics across the entire heart. Therefore, body surface potential (BSP) mapping has been developed by deploying a larger number of ECG electrodes, such as Dalhousie Standard Torson 120 electrodes, to provide high-resolution spatiotemporal information about cardiac electrodynamics projected onto the body surface [8]. It allows for detailed mapping and localization of cardiac activity and thus has been widely used in the last decades for solving the ECG inverse problem. Nevertheless, the ECG inverse problem, like many other inverse problems, faces two primary challenges. Firstly, the problem is ill-posed, meaning that it may not have a unique solution and small errors in the input data can lead to significant errors in the output. Secondly, solving the ECG inverse problem can be computationally intensive, particularly when dealing with high-resolution BSP mapping data. The complex geometrical, mechanical, and electrical properties of the heart, as well as the need for real-time or near-real-time solutions in clinical applications, pose computational challenges.
Figure 1: Illustration of the electrocardiogram (ECG) inverse problem.
@@
Term “Electrocardiogram” or “ECG” or “EKG” or “electrocardiographic” or “electrophysiological” or “electrophysiology” or “EP” or “electrocardiography” or “body surface potential mapping” or “BSPM” and
@@
“Inverse problem/ issue” or “inverse solution/ estimation/ inference/ reconstruction” or “cardiac model personalization” or “electrocardiographic imaging/ mapping” or “ECG-imaging” or “ECGI” or “activation imaging” or “noninvasive imaging” and
@@
“Electrical activity/ excitation” or “surface potential” or “epicardial/ endocardial potential” or “action potential” or “transmembrane potential” or “activation sequence” or “activation wavefront” or “activation time map” or “pacing site” or “atrial potential” or “local parameter”
Figure 2: Timeline of the development of ECG, imaging, and cardiac digital twins.
1.1 History: From Electrocardiography to Cardiac Digital Twins
Investigations of cardiac electrophysiology can be traced back to 1842 when Carlo Matteucci identified signs of cardiac electrical activity in frog hearts [9]. Dr. Augustus Waller furthered this research by recording the first human electrocardiogram in 1887, using a capillary electrometer and electrodes placed on the chest and back of a human subject [10]. His study demonstrated that electrical activity preceded ventricular contraction. Subsequently, in 1893, Einthoven coined the term “electrocardiogram” to describe the cardiac waveforms [11]. The American Heart Association (AHA) standardized the 12-lead ECG in 1954 [12], paving the way for advancements in cardiac electrophysiology. In 1960, Denis Noble developed one of the first mathematical models of a cardiac electric cell based on Hodgkin-Huxley equations, further refining it to characterize the long-lasting action and pacemaker potentials observed in the Purkinje fibres of the heart [13, 14, 15]. This laid the groundwork for cardiac simulations using computer models. Meanwhile, body surface potential mapping (BSPM) emerged as an alternative method for non-invasive and high-resolution measurement of human cardiac electrical activity in 1963 [16]. In 1970, Durrer et al. [17] conducted a seminal investigation into the electrical behavior of the human heart at the organ level. Since 1972, numerous studies have tackled the electrocardiography inverse problem using analytical and computational models [18, 19, 20, 21], experimental animals [22, 23], and human subjects [24, 25]. Advancements in medical imaging have also enabled non-invasive visualization of cardiac anatomy since the inception of techniques like the first human MRI scan in 1977 [26]. The formalization of inverse procedures reconstructing bioelectric sources from BSPM and medical imaging led to the development of electrocardiographic imaging (ECGi) techniques, notably pioneered by Yoram Rudy’s lab [27, 28, 29, 30, 31].
@@ 1.1 History: From Electrocardiography to Cardiac Digital Twins
The concept of the digital twin was formally introduced in 2002 by Michael Grieves, though it can be traced back to NASA’s use for the Apollo 13 rescue mission in 1970 [32]. The Physiome Project was initiated in 2003 with the aim of establishing a framework for modelling the human body, utilizing computational methods to integrate biochemical, biophysical, and anatomical information across cells, tissues, and organs [33]. Starting from 2014, many companies like Dassault, Siemens, and ANSYS embraced the term “digital twins” in their marketing initiatives. For example, Dassault Systèmes initiated the Living Heart Project in 2014, which is the first comprehensive computer model of the human heart, integrating all functional aspects, i.e., blood flow dynamics, mechanical properties, and electrical conduction [34]. In recent years, digital twin tools have become popular in healthcare, such as cardiac digital twins [35], brain digital twins [36], etc. In 2016, the US Food & Drug Administration (FDA) Center for Devices and Radiological Health released the first guidance to enable the use of modelling and simulation to develop in silico trials [37]. In 2019, Trayanova’s lab received FDA approval for a randomized clinical trial involving 160 patients, termed OPTIMA (OPtimal Target Identification via Modeling of Arrhythmogenesis), to demonstrate the effectiveness of cardiac digital twins in guiding atrial ablation procedures [3]. The brief timeline has been presented in Fig. 2.
1.2 Study Inclusion and Literature Search
In this work, we aim to provide readers with a survey of the state-of-the-art solutions to electrocardiographic inverse problems. There exist several modalities for cardiac electrical activity assessment, such as ECG, BSP mapping, electrograms (EGM), electro-anatomic mapping (EAM), vectorcardiography (VCG), etc. However, in the context of solving the ECG inverse problem, we usually only involve ECG and BSP as the observed data. Note that compared to standard 12-lead ECG, BSP mapping is, however, not yet widely available as a diagnostic modality [38]. For cardiac activity representation, two major source models are employed, i.e., heart surface potential (HSP) and transmembrance potential defined on the volumetric mesh of the heart. Specifically, it can be further categorized into epicardial and/ or endocardial potential, activation time map/ activation wavefront, transmembrane potential, action potential duration map, voltage potential maps, isochrones/ isochrone maps, etc.
@@ 1.2 Study Inclusion and Literature Search
Table 2: Summary of the related representative existing surveys. DL: deep learning; MCG: Magnetocardiography; ATM: activation time maps; AF: atrial fibrillation.
Source & Publish Year Scope Inclusion
Macleod et al. (1998) [7] ECG inverse problem conventional solutions
Dossel et al. (2000) [39] ECG and MCG inverse problem conventional solutions
Yaman et al. (2013) [40] inverse problem for applied sciences not for ECG; conventional solutions
@@
Arridge et al. (2019) [44] inverse problem not for ECG; mainly DL-based solutions
Sameni et al. (2010) [45] fetal ECG analysis fetal ECG; partially includes inverse problem
Bear et al. (2015) [46] ECG forward problem not for inverse problem
Hong et al. (2020) [47] DL-based ECG analysis not for inverse problem
Liu et al. (2021) [48] DL-based ECG analysis not for inverse problem
@@
Bifulco et al. (2021) [49] translational applications of computational modelling translational applications
Cantwell et al. (2015) [50] ECG inverse problem mainly focusing on the inference of ATM and CV
Cluitmans et al. (2015) [51] ECG inverse problem mainly focusing on conventional methods
Cluitmans et al. (2018) [52] ECG inverse problem mainly focusing on the validation and clinical application
Yadan et al. (2023) [53] ECG inverse problem mainly focusing on the identification of AF drivers
Hernandez et al. (2023) [54] ECG inverse problem mainly focusing on the atria
To ensure comprehensive coverage, we have screened publications mainly from the last 24 years (2000-2024) related to this topic. Our main sources of reference were Internet searches using engines, including Google Scholar, PubMed, IEEE-Xplore, Connected Papers, and Citeseer. To cover as many related studies as possible, flexible search terms have been employed when using these search engines, as summarized in Table 1. There are several research teams and industry companies at the forefront of computational cardiology and electrocardiography, working to advance our understanding of cardiac function, arrhythmias, and heart-related disorders through computational methods and simulations  [^1] Continuously updating compilation of cardiac digital twin resources: github.com/lileitech/Awesome-Cardiac-Digital-Twins. We followed their recent works to obtain state-of-the-art methodologies and clinical applications. Both peer-reviewed journal papers and conference papers were included here. In the way described above, we have collected a comprehensive library of more than 200 papers. Note that we picked the most detailed and representative ones for this review when we encountered several papers from the same authors about the same subject.
@@ 1.3 Existing Survey from Literature
Table 2 lists current review papers related to the topic, i.e., ECG inverse problem. One can see that the scopes of current review works are different from ours, though with partial overlaps. For example, there are several existing reviews focusing on summarizing solutions of inverse problems but not for ECG data [40, 41, 42, 44, 43]. Alternatively, there exist few surveys on the ECG analysis, but not focusing on the inverse problem [46, 48, 1]. Bifulco et al. [49] provided the potential clinical applications of the ECG inverse inference strategies, but it is not a methodological survey. Although Macleod et al. [7] and Dossel et al. [39] performed a methodology survey on the ECG inverse problem, we focus on the developments proceeding their reviews. Cantwell et al. [50] mainly summarized the techniques for local activation time annotation and conduction velocity estimation in cardiac mapping. We provide a comprehensive and up to date review of ECG inverse inference, encompassing both its recent methodological aspects and potential clinical applications.
@@ 2.1 Definition of the Problem
The ECG inverse problem refers to the mathematical task of determining the electrical activity within the heart based on measurements of the ECG signals recorded on the body surface. It involves applying a quasi-static approximation of Maxwell’s equations on the electromagnetic field between the heart and the torso to establish a relationship between the ECG data and the cardiac electrical excitation. Solving these equations numerically on subject-specific heart-torso models yields a linear measurement model,
@@
where $\boldsymbol{y}(\boldsymbol{s},t)$ and $\boldsymbol{u}(\boldsymbol{s},t)$ refer to the vector of ECG signals and cardiac electrical parameters at position $\boldsymbol{s}$ and time instant $t$, $\boldsymbol{H}$ denotes the forward matrix, and $\epsilon$ represents the noise term that accounts for measurement and modelling errors. Based on this physical system modelled by PDEs and/or ordinary differential equations (ODEs), the typical solution of the problem is to inversely estimate $\boldsymbol{u}(\boldsymbol{s},t)$ that minimizes the fitting of the observational data $\boldsymbol{y}(\boldsymbol{s},t)$,
$\min_{\boldsymbol{u}(\boldsymbol{s},t)}\|\boldsymbol{y}(\boldsymbol{s},t)-\boldsymbol{H}\boldsymbol{u}(\boldsymbol{s},t)\|_{2}^{2},$ (2)
which however is ill-posed. The inverse problem can exhibit one of three characteristics: it may have multiple solutions (non-unique), no solution (non-existence), or solutions that are highly sensitive to small perturbations in the measured data (unstable) [55]. Moreover, it is typically influenced by measured input data that is frequently corrupted by random or stochastic noise.
To solve the ECG inverse problem, the state-of-the-art approaches can be coarsely separated into two kinds: deterministic and probabilistic methods [56], as presented in Fig. 3. Deterministic approaches in cardiac electrophysiology involve minimizing a cost function that quantifies the discrepancy between the observed data and the model predictions. Rather than aiming solely to minimize a cost function, probabilistic models embrace uncertainty and provide a probabilistic distribution over possible outcomes.

### v4

@@ Abstract
Cardiac digital twins (CDTs) are personalized virtual representations used to understand complex cardiac mechanisms. A critical component of CDT development is solving the ECG inverse problem, which enables the reconstruction of cardiac sources and the estimation of patient-specific electrophysiology (EP) parameters from surface ECG data. Despite challenges from complex cardiac anatomy, noisy ECG data, and the ill-posed nature of the inverse problem, recent advances in computational methods have greatly improved the accuracy and efficiency of ECG inverse inference, strengthening the fidelity of CDTs. This paper aims to provide a comprehensive review of the methods of solving ECG inverse problem, the validation strategies, the clinical applications, and future perspectives. For the methodologies, we broadly classify state-of-the-art approaches into two categories: deterministic and probabilistic methods, including both conventional and deep learning-based techniques. Integrating physics laws with deep learning models holds promise, but challenges such as capturing dynamic electrophysiology accurately, accessing accurate domain knowledge, and quantifying prediction uncertainty persist. Integrating models into clinical workflows while ensuring interpretability and usability for healthcare professionals is essential. Overcoming these challenges will drive further research in CDTs.
@@
Cardiac digital twins, ECG, inverse problem, ECGI, parameter estimation, model personalization, survey.
@@ 1 Introduction
Electrocardiography (ECG) investigates the relationship between the cardiac electrical activity and the resulting potential measured on the torso surface. Understanding this relationship is crucial for developing personalized virtual models of cardiac electrophysiology (EP), i.e., cardiac digital twins (CDTs) [1, 2]. Currently, treatment decisions are still primarily based on empirical clinical studies that statistically compare the effects of different treatments across large groups of patients with similar cardiac conditions [3]. Instead, the virtual heart models have shown great promise in providing personalized treatments, such as arrhythmia risk stratification [4] and ablation guidance for persistent atrial fibrillation (AF) [5]. This involves using computational models to solve the ECG forward and inverse problems. The ECG forward problem aims to predict signal on the body surface from known cardiac activity. Instead, the ECG inverse problem seeks to infer cardiac information from ECG or body surface potential (BSP) [6, 2], as presented in Fig. 1. It can be conducted in two aspects, i.e., the reconstruction of cardiac sources and the estimation of EP parameters. Both aspects are closely linked with the concept of CDTs [7, 2]. The reconstruction of cardiac sources provides crucial information about the cardiac electrical activity, which can be integrated into digital twin models to enhance their accuracy in representing the heart [7]. On the other hand, the estimation of EP model parameters allows these digital twins to be fine-tuned to reflect individual patient conditions, ensuring that the model behavior closely matches the real cardiac physiological responses [7].
Figure 1: Illustration of the electrocardiogram (ECG) inverse problem, including non-invasive cardiac source reconstruction, also known as ECG imaging (ECGI), and electrophysiological (EP) parameter estimation. Here, we take the biventricular modeling as an example. BSP: body surface potential; CV: conduction velocity.
However, due to the nature of the ECG inverse problem, the reconstruction of cardiac sources and the estimation of EP parameters share several common challenges. First, both scenarios involve solving an inherently ill-posed inverse problem, which may not have a unique solution. This means that small errors in the input data can lead to significant inaccuracies in the output. Second, ECG data recorded from the body surface, often using a limited number of electrodes such as the 12-lead ECG with ten electrodes, has limited spatial resolution and is susceptible to noise, artifacts, and inconsistencies. Third, both scenarios need to account for the complex, inhomogeneous structure of the thorax, including the heart, lungs, and other tissues, which influence how electrical signals propagate to the body surface. Accurately modeling these structures poses a significant challenge that affects the reliability of reconstructed cardiac sources and the accuracy of estimated EP parameters. Compared to the inverse inference of cardiac sources, the inference of EP parameters presents several additional challenges. While the cardiac source reconstruction requires computational resources, it is generally not considered excessively time-consuming. In contrast, EP parameter estimation is far more computationally intensive due to its reliance on sophisticated anatomical modeling, iterative simulations, and complex parameter fitting, particularly when using high-resolution models. Also, EP parameters can be interdependent, meaning that changes to one parameter may affect others, which further complicates the optimization and fitting process [8]. Unlike standard ECG imaging, which is directly validated against real-time measurements, the EP parameter estimation requires comprehensive validation to ensure the generated CDTs accurately represent the unique anatomy and EP of the patient.
1.1 History: From Electrocardiography to Cardiac Digital Twins
Investigations of cardiac electrophysiology can be traced back to 1842 when Carlo Matteucci identified signs of cardiac electrical activity in frog hearts [9]. Dr. Augustus Waller furthered this research by recording the first human electrocardiogram in 1887, using a capillary electrometer and electrodes placed on the chest and back of a human subject [10]. His study demonstrated that electrical activity preceded ventricular contraction. Subsequently, in 1893, Einthoven coined the term “electrocardiogram” to describe the cardiac waveforms [11]. The American Heart Association (AHA) standardized the 12-lead ECG in 1954 [12], paving the way for advancements in cardiac electrophysiology. In 1960, Denis Noble developed one of the first mathematical models of a cardiac electric cell based on Hodgkin-Huxley equations, further refining it to characterize the long-lasting action and pacemaker potentials observed in the Purkinje fibres of the heart [13, 14, 15]. This laid the groundwork for cardiac simulations using computer models. Meanwhile, BSP mapping emerged as an alternative method for non-invasive and high-resolution measurement of human cardiac electrical activity in 1963 [16]. In 1970, Durrer et al. [17] conducted a seminal investigation into the electrical behavior of the human heart at the organ level. Since 1972, numerous studies have tackled the electrocardiography inverse problem using analytical and computational models [18, 19, 20, 21, 22], experimental animals [23, 24], and human subjects [25, 26]. The formalization of inverse procedures reconstructing bioelectric sources from BSP mapping (BSPM) led to the development of non-invasive electrocardiographic imaging (ECGI), a term formally coined by Yoram Rudy’s lab in 1997 [27]. While ECGI can be patient-specific in terms of geometry and electrical potentials, it does not typically include a full-scale, personalized model of the cardiac EP properties.
@@ 1.1 History: From Electrocardiography to Cardiac Digital Twins
Term “Electrocardio*” or “ECG” or “EKG” or “electrophysiolog*” or “EP” or “body surface potential” or “BSP” and
@@
“Inverse problem/ issue” or “inverse solution/ estimat*/ infer*/ reconst*” or “electrocardiographic imaging/ mapping” or “ECG-imaging” or “ECGI” or “activation imaging” or “noninvasive imaging” or “personaliz*/ calibrat*” and
@@
“Electrical activity/ excitation” or “heart/ epicardial/ endocardial /atrial / transmembrane potential” or “activation sequence/ wavefront/ time” or “pacing site” or “parameter” or “onset location/ root note/ earliest activation site/ Purkinje” or “conducti*” or “origin of cardiac activation” or “excitability”
@@
In constrast, CDT is a comprehensive, personalized model that incorporates both the electrical and structural properties of the heart, including detailed EP parameters. The concept of the digital twin was formally introduced in 2002 by Michael Grieves, though it can be traced back to NASA’s use for the Apollo 13 rescue mission in 1970 [28]. The Physiome Project was initiated in 2003 with the aim of establishing a framework for modeling the human body, utilizing computational methods to integrate biochemical, biophysical, and anatomical information across cells, tissues, and organs [29]. Starting from 2014, many companies like Dassault, Siemens, and ANSYS embraced the term “digital twins” in their marketing initiatives. For example, Dassault Systèmes initiated the Living Heart Project in 2014, which is the first comprehensive computer model of the human heart, integrating many functional aspects, i.e., blood flow dynamics, mechanical properties, and electrical conduction [30]. In recent years, digital twin tools have become popular in healthcare, such as CDTs [31], brain digital twins [32], etc. In 2016, the US Food & Drug Administration (FDA) Center for Devices and Radiological Health released the first guidance to enable the use of modeling and simulation to develop in silico trials [33]. Notably, Trayanova’s lab received FDA approval in 2019 for a randomized clinical trial involving 160 patients, termed OPTIMA (OPtimal Target Identification via Modeling of Arrhythmogenesis), to demonstrate the effectiveness of CDTs in guiding atrial ablation procedures [5].
@@
Table 2: Summary of the related representative existing surveys. DL: deep learning.
Source & Publish Year Scope Gap
Corral et al. (2020) [31] cardiac digital twins challenges and clinical applications
Coorey et al. (2022) [34] cardiac digital twins challenges and clinical applications
Bear et al. (2015) [35] ECG forward problem not for inverse problem
Bergquist et al. (2021) [36] ECG forward problem not for inverse problem
Gulrajani et al. (1998) [37] ECG forward and inverse problem conventional solutions
@@
Siontis et al. (2021) [41] DL-based ECG analysis not for inverse problem
@@
Bifulco et al. (2021) [43] ECG inverse problem applications of computational modeling
Macleod et al. (1998) [6] ECG inverse problem conventional ECGI solutions
Cluitmans et al. (2018) [7] ECG inverse problem conventional ECGI solutions and applications
Hernandez et al. (2023) [44] ECG inverse problem atrial ECGI
Loewe et al. (2022) [45] ECG inverse problem partially EP parameter estimation
1.2 Study Inclusion and Literature Search
In this work, we aim to provide readers with a survey of the state-of-the-art solutions to ECG inverse problems. There exist several modalities for cardiac electrical activity assessment, such as ECG, BSP, electrogram (EGM), electro-anatomic mapping (EAM), vectorcardiography (VCG). However, in the context of solving the ECG inverse problem, we usually only involve non-invasive ECG and BSP as the observed data. Note that compared to standard 12-lead ECG, BSPM is, however, not yet widely available as a diagnostic modality [46]. For cardiac activity representation, three major source models are employed, i.e., surface potential models, activation models, and transmembrane potential (TMP). For model personalization, the EP parameters to be estimated can be conductivities, conduction velocities (CVs), earliest activation sites (EASs), tissue excitability, etc.
To ensure comprehensive coverage, we have screened publications mainly from the last 24 years (2000-2024) related to this topic. Our main sources of reference were Internet searches using engines, including Google Scholar, PubMed, IEEE-Xplore, Connected Papers, and Citeseer. To cover as many related studies as possible, flexible search terms have been employed when using these search engines, as summarized in Table 1. There are several research teams and industry companies at the forefront of computational cardiology and electrocardiography, working to advance our understanding of cardiac function, arrhythmias, and heart-related disorders through computational methods and simulations  [^1] Continuously updating compilation of cardiac digital twin resources: github.com/lileitech/Awesome-Cardiac-Digital-Twins. We followed their recent works to obtain state-of-the-art methodologies and clinical applications. Both peer-reviewed journal and conference papers were included here. In the way described above, we have collected a comprehensive library of more than 200 papers. Note that we normally only picked the most detailed and representative ones for this review when we encountered several papers from the same authors about the same subject.
@@ 1.3 Existing Survey from Literature
Table 2 lists current review papers related to the topic, i.e., ECG inverse problem. One can see that the scopes of current review works are different from ours, though with partial overlaps. For example, several surveys have been conducted to summarize the current challenges and existing or potential clinical applications of ECGI [7] or CDTs [31, 34]. Furthermore, there are several existing reviews focusing on summarizing solutions of inverse problems but not for ECG data [38, 39, 40]. Alternatively, there exist few surveys on the ECG analysis, but not focusing on the inverse problem [41, 42]. Bifulco et al. [43] provided the potential clinical applications of the ECG inverse inference strategies, but it is not a methodological survey. Although Macleod et al. [6], Cluitmans et al. [47] and Hernandez et al. [44] performed a methodology survey on the ECG inverse problem, they only summarized the ECG inverse inference studies for cardiac source reconstruction, i.e., ECGI. Loewe et al. [45] partially reviewed the ECG inverse inference works on EP parameter estimation in terms of cellular and tissue levels. Here, we provide a comprehensive and up-to-date review of ECG inverse inference in a broader range, encompassing both cardiac source reconstruction and EP parameter estimation. The communities engaged in these two areas of research do not completely overlap and sometimes even use different terminologies. For example, in the narrow definition of the inverse-ECG problem, i.e., ECGI, the forward model typically refers to the relationship between cardiac sources and ECG data. In contrast, within the model personalization community, the forward model often pertains to the forward EP model. This survey aims to bridge the gap between these communities, fostering conversation and collaboration.
@@ 2.1 Definition of the Problem
The ECG inverse problem involves inferring internal cardiac information from surface ECG measurements. This problem can be formulated in two ways:
  1. 1.
    Reconstruction of cardiac sources (ECGI): It aims to reconstruct the cardiac electrical activity from the recorded ECG or body surface potentials (BSPs) for a non-invasive visualization of cardiac activity. The relationship between the ECG data $\boldsymbol{y}(\boldsymbol{s},t)$ and the cardiac sources $\boldsymbol{u}(\boldsymbol{s},t)$ at position $\boldsymbol{s}$ and time instant $t$ is typically modeled as a linear system:
@@
    where $\boldsymbol{H}$ is approximated as a forward matrix, and $\epsilon$ accounts for noise. The goal is to estimate $\boldsymbol{u}(\boldsymbol{s},t)$ by minimizing the fitting of the observational data $\boldsymbol{y}(\boldsymbol{s},t)$:
    $\min_{\boldsymbol{u}(\boldsymbol{s},t)}\|\boldsymbol{y}(\boldsymbol{s},t)-\boldsymbol{H}\boldsymbol{u}(\boldsymbol{s},t)\|_{2}^{2}.$ (2)
    Note that this linear model specifically applies to potential-based ECG inverse problems. Activation-based models, which often involve non-linear relationships, are not directly addressed by this formulation [48, 49, 50].
  2. 2.
    Estimation of EP model parameters: It focuses on personalize cardiac models by estimating patient-specific EP parameters within a mathematical model of cardiac electrophysiology, such as the Aliev-Panfilov (AP) model [51] or Eikonal model [1]. The AP model can be defined as,
    $\frac{\partial u}{\partial t}$ $=\nabla\cdot(\boldsymbol{D}\nabla\boldsymbol{D})-cu(u-\theta)(u-1)-\frac{\partial u}{\partial z},$ (3)
    $\frac{\partial z}{\partial t}$ $=\epsilon(u,z)\left(-z-cu(u-\theta-1)\right),$
    where $u$ is the normalized transmembrane action potential, $z$ is the recovery current, $\boldsymbol{D}$ is the diffusion tensor, $c$ controls the repolarization, and $\theta$ is tissue excitability parameter. The Eikonal equation describes the propagation of the activation wavefront:
    $\begin{cases}|\nabla\tau(\boldsymbol{s})|=\dfrac{1}{v(\boldsymbol{s})},&\boldsymbol{s}\in\Omega,\\ \tau(\boldsymbol{s})=\tau_{0}(\boldsymbol{s}),&\boldsymbol{s}\in\Gamma_{0},\end{cases}$ (4)
    where $\tau(\boldsymbol{s})$ represents the activation time at point $\boldsymbol{s}$ within the cardiac tissue domain $\Omega$, while $v(\boldsymbol{s})$ denotes the local CV. The boundary conditions are defined by the set of EASs $\Gamma_{0}$, with $\tau_{0}(\boldsymbol{s})$ specifying the activation times at these sites. Taking the Eikonal model as an example, its parameter optimization objective function can be formulated as,
    $\min_{\boldsymbol{v},\tau_{0}}\|\boldsymbol{y}_{\text{ECG}}-\boldsymbol{y}_{\text{model}}(\boldsymbol{v},\tau_{0})\|_{2}^{2}+\lambda\mathcal{R}(\boldsymbol{v}),$ (5)
    where $\boldsymbol{y}_{\text{ECG}}$ is the measured signal, while $\boldsymbol{y}_{\text{model}}(\boldsymbol{v},\tau_{0})$ represents the simulated ECG signals using the forward EP model with CVs $\boldsymbol{v}$ and EASs $\tau_{0}$; $\mathcal{R}(\boldsymbol{v})$ imposes certain constraints on the estimated parameters, and $\lambda$ is a regularization term.
Both are essential for creating CDTs, but they may involve different methods. The state-of-the-art approaches can be coarsely separated into two kinds: deterministic and probabilistic methods for both [52], as presented in Fig. 2. Deterministic approaches in cardiac electrophysiology involve minimizing a cost function that quantifies the discrepancy between the observed data and the model predictions. Rather than aiming solely to minimize a cost function, probabilistic models embrace uncertainty and provide a probabilistic distribution over possible outcomes. We will review these two approaches separately and summarize their application to different formulations of ECG inverse inference. While we categorize the methodologies into deterministic and probabilistic, it is important to note that these categories often overlap, which will be emphasized when introducing these methods.